OpenStudy (whatever):

sometimes it is easier to use 2 variables when translating a word problem. then you must find a system of 2 equations to solve in order to find the solution to the problem. translate to a system of equation and solve. roberta is 25 years older than her daughter, cindy. in 4 years, roberta will be twice as old as cindy. how old are they now?

OpenStudy (anonymous):

r=c+25 r=2c+4

OpenStudy (anonymous):

sorry r=2c-4

OpenStudy (whatever):

how does that equation work?

OpenStudy (anonymous):

i just read the data...

OpenStudy (anonymous):

let roberta's present age be x and let cindy's present age be y x=25+y after four years, roberts's age will be x+4 and cindy's age will be y+4 x+4=2(y+4)

OpenStudy (anonymous):

now you can substitute x=25+y in x+4=2y+8

OpenStudy (whatever):

does y equal 2 ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

25+y+4=2y+8 29-8=2y-y y=21

OpenStudy (anonymous):

so x=25+y x=25+21 x=46

OpenStudy (anonymous):

so the present age of roberta is 46 and cindy is 21

OpenStudy (whatever):

ohh now i get it

OpenStudy (anonymous):

if robert's age = "x", & cindy's age ="y" at present, x =25+y and y = x - 25, therefore, if y = 0, x = 25.

OpenStudy (anonymous):

therefore in 4 yrs, x = 2y+4 so, if x = y + 25 ---- (1) x = 2y +4 ----(2) substitute (1) in 2 y+25 = 2y+4 y = 21 therefore at present, cindi = 21, and robert is 46

OpenStudy (whatever):

thank you